OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? A: y = 2x + 1 B: y = 2x + 3 C: y = 2x - 8 D: y = 2x - 3

OpenStudy (anonymous):

use equation of line y=mx+c where m is slope substitute slope then substitute x and y values to get c

OpenStudy (anonymous):

i got y = 2x + 2/5 ? and that's not right

OpenStudy (anonymous):

it is a

OpenStudy (anonymous):

y=mx+c 5=2(2)+c c=1

OpenStudy (anonymous):

y=2x+1

OpenStudy (anonymous):

\[y-y _{1}=m (x-x _{1})\] y-5 =2(x-2) y-5 =2x-4 y = 2x-4+5 y=2x+1

OpenStudy (anonymous):

Thank you sooo much! all of youu.